Mary works due north of home. her husband juan works due east of home.they leave for work at same time. by the time mary is 7 miles from home the distance between them is 1 mile more than the distance juan is from home. how far from home is juan?
Answer by cleomenius(959)

What we will be doing here is to set a right triangle, the legs will be 7 and x, and the hypotneus is x +1. This is the present distance between Mary and Juan.
x^2 + 49 = (x + 1)^2
x^2 + 49 = x^2 + 2x + 1
48 = 2x
24 = x
7^2 + 24^2 = 25^2 to check.
49 + 576 = 625.
This does check.
Cleomenius.